OpenStudy (anonymous):

2x/5-2/3=x/2 Please can you show it step by step as well

OpenStudy (anonymous):

first you have to find the common denominator

OpenStudy (anonymous):

Use a common denomiator, 15 6x/15 - 10/15 = 7.5x/15 (6x-10)/15 = 7.5x/15 6x-10 = 7.5x -10 = 1.5x x = -20/3

OpenStudy (anonymous):

2x*3/5*3-2*5/3*5=x/2 6x/15-10/15=x/2 6x-10=15x/2 12x-20=15x -20=15x-12x -20=3x -20/3=x